How do I find different ways to write log(100)using logs and powers using only intergers 0, 1, 2, 3, 10 and logs?
Answer by venugopalramana(3286)

How do I find different ways to write log(100)using logs and powers using only intergers 0, 1, 2, 3, 10 and logs?
EXPRESS 100 AS A COMBINATION OF PRODUCT OR QUOTIENT OF GIVEN NUMBERS AND USE
LOG(X*Y)=LOG(X)+LOG(Y)
LOG(X/Y) = LOG(X) - LOG(Y)
EX... 100= 2*5*10
HENCE LOG (100) = LOG(2*5*10) = LOG(2)+ LOG(5) + LOG(10) ETC...